Give an example, with justification, of a 

skew-Hermitian operator on C2.



1
Expert's answer
2021-07-12T07:14:24-0400


Let there be a matrix A



A = "\begin{bmatrix}
 i &1-i & 2\\
 -1-i & 3i & i\\
-2& i & 0
\end{bmatrix}"


Now A is skew Hermitian only if



( A* )T = - A



Now we take the Conjugate of A and that is given as


A"\begin{bmatrix}
 -i & 1+i& 2\\
 -1+i & -3i & -i\\
-2&-i&0
\end{bmatrix}"




On taking the transpose of above we have



( A)T = "\begin{bmatrix}
 -i & -1+i & -2\\
 1+i &-3i & -i\\
2&-i&0
\end{bmatrix}"




Now,


- A = "\begin{bmatrix}
 -i & -1+i & -2\\
 1+i &-3i & -i\\
2&-i&0
\end{bmatrix}"




On comparing ( A* )T and - A we see that


( A* )T = - A


Hence, A is a skew Hermitian matrix.